### v1

# Learning from Heterogeneity: A Dynamic Learning Framework for Hypergraphs

Abstract
Graph neural network (GNN) has gained increasing popularity in recent years owing to its capability and flexibility in modeling complex graph structure data. Among all graph learning methods, hypergraph learning is a technique for exploring the implicit higher-order correlations when training the embedding space of the graph. In this paper, we propose a hypergraph learning framework named Learning from Heterogeneity (LFH) that is capable of dynamic hyperedge construction and attentive embedding update utilizing the heterogeneity attributes of the graph. Specifically, in our framework, the high-quality features are first generated by the pairwise fusion strategy that utilizes explicit graph structure information when generating initial node embedding. Afterwards, a hypergraph is constructed through the dynamic grouping of implicit hyperedges, followed by the type-specific hypergraph learning process. To evaluate the effectiveness of our proposed framework, we conduct comprehensive experiments on several popular datasets with eleven state-of-the-art models on both node classification and link prediction tasks, which fall into categories of homogeneous pairwise graph learning, heterogeneous pairwise graph learning, and hypergraph learning. The experiment results demonstrate a significant performance gain (average 12.5% in node classification and 13.3% in link prediction) compared with recent state-of-the-art methods.
Index Terms:
Heterogeneous Hypergraph, Representation Learning, Hypergraph Generation, Classification.
I Introduction
Graph learning has attracted tremendous attention in recent years owing to its prominent capability when modelling structure-based data. In particular, there has been witnessed an increasing use of graph models (i.e., GNN: Graph Neural Network) in many fields of applications, such as social network recommendation [1], medical diagnosis [2, 3], and text analysis [4]. The goal of graph learning is to encode the graph structure of different input data into an embedding space, where the representations can be used for downstream node/edge/graph tasks.
There are both standard graphs and hypergraphs. Fig. 1 illustrates (a) a standard graph where one edge connects two vertices regardless of the edge type and (b) a hypergraph with three hyperedges. Standard graphs can be modeled using many existing GNN models such as graph convolution network (GCN) [5], graph attention network (GAT) [6], and heterogeneous graph learning (PC-HGN) [7]. Most of these models aim to model the pairwise relationships in well-structured graph data. However, this may not work for hypergraphs, as a hypergraph can encompass any number of vertices (nodes) at each edge type, presenting higher-order correlations with rich underlying information. Pairwise graph learning models, in this case, become intractable. Specifically, hypergraph presents more complex non-pairwise relationships, where these implicit relationships are not only dyadic (pairwise) but rather triadic or even higher [8]. In general, hypergraph learning can be regarded as a generalized form of traditional graph learning.
Like pairwise graph learning models like GCN, hypergraph learning can also be seen as a process of uncovering and passing the structure information among different hyperedges, during which low-dimensional node/graph level embedding representations can be learned and used for downstream tasks, such as node classification [9], link prediction [10], and clustering [11].
Figure 1: Comparison between (a) a standard graph with 5 nodes and 4 edges, and (b) a hypergraph with 3 hyperedges (the three colored node groups).
Owing to the great potential of exploring high-order correlations among the data, hypergraph learning has drawn increasing attention from the academic community and has been employed in applications such as image classification [12], video segmentation [13], and hyperspectral image analysis [14]. How to construct hyperedges lies at the core of hypergraph learning. Existing works generally take two approaches, i.e., attribute-based construction and neighborhood-based construction. The early work focused on constructing the hypergraph based on the features of each vertex [15]. This strategy instructs the model to use only the feature information which inevitably sacrifices certain generality from the topological perspective. The neighborhood-based strategy addresses this weakness by enabling the locality concept. Specifically, a hyperedge is constructed through clustering a master vertex and its neighbor vertices [16]. However, the clustering method (e.g., k-nearest neighbors (k-NN)) separates hypergraph construction from the graph learning process, and thus is sensitive to noisy data, especially in visual classification tasks. Moreover, most neighborhood-based models fail to incorporate topological heterogeneity into the learning process, which limits the learning to a single type of hyperedges.
To tackle these challenges, we propose a dynamic learning framework, Learning from Heterogeneity (LFH), to improve the quality of representation learning for hypergraphs. Our LFH framework is composed of three key modules: 1) initial embedding generation, 2) dynamic hypergraph construction, and 3) attention-based heterogeneous hypergraph learning. Concretely, initial embedding generation aims to exploit the pairwise connectivities in the graph, helping to fuse the explicit topological information into the initial embedding space. Following that, as an integral part of representation learning, the heterogeneous hypergraph is constructed and then learned dynamically.
In summary, the main contributions of this work are:
  • We propose a novel hypergraph learning framework LFH with three modules: initial embedding generation, dynamic hypergraph construction, and attention-based heterogeneous hypergraph learning. A pairwise fusion strategy is also considered to fully exploit the explicit pairwise graph information to generate high-quality initial embedding.
  • We design a dynamic learning process to generate different types of implicit hyperedges to construct the hypergraph and dynamically adapt the hypergraph construction during the learning process. The embedding is updated iteratively through a type-specific multi-head attention mechanism, so as to encode the heterogeneous attributes into the embedding space.
  • We conduct extensive experiments on two different downstream tasks, including node classification and linked prediction, and show that our LFH outperforms state-of-the-art homogeneous pairwise graph learning models, heterogeneous pairwise graph learning models, and hypergraph learning models consistently across all datasets by a large margin.
The remainder of this paper is organized as follows. Section 2 reviews the related works of graph learning on homogenous, heterogenous, and hypergraphs. Section 3 provides the preliminary knowledge of hypergraphs, while Section 4 introduces the detail of our proposed LFH framework, along with analyses and discussions. The experimental results are reported and analyzed in Section 5. Finally, Section 6 concludes the paper.
II Related Work
II-A Homogeneous Graph Learning
As a basic graph structure, a homogeneous graph consists of a single node type with a single relation. Typical GNN models for homogeneous graphs include GCN [5] and GAT [6]. GCN extends the traditional convolutional neural networks (CNN) to handle graph-structured data by performing convolutions in the spectral/spatial domain of the graph nodes to capture the structural information of the graph. On the other hand, GAT applies the self-attention mechanism to assign dynamic weights for the neighbors of a node and then take a weighted sum of their embeddings to obtain the node’s representation. This enables GAT to learn adaptive neighborhood representations and capture complex relationships between nodes.
The primary difference between various GNN models lies in the way how messages are passed between the nodes to learn the representation. Instead of training individual embeddings for each node, GraphSAGE [17] leverages the node features in the learning algorithm to train a set of aggregator functions to generate embeddings for entirely unseen nodes. Aiming for visual question answering (VQA) services, Li et al. proposed a GAT-based platform that encodes each image into a graph and models multi-type inter-object relations to learn relation representation from the graphs [18]. Jiang et al. presented a GCN-based framework (GLCN) for graph data representation learning. GLCN generates similarity-based graph structure by simultaneous graph learning and graph convolution in a unified network architecture [19]. Compared with the embedding frameworks that can only generate embeddings for a single fixed graph like transductive learning in GCN, Zeng et al. proposed GraphSAINT, a graph sampling based inductive representation learning method, to generalize across different graphs. GraphSAINT generates low-dimensional vector representations for the nodes and is beneficial for large graphs with rich node attribute information [20].
II-B Heterogeneous Graph Learning
Different from homogeneous graphs, the nodes in a heterogeneous graph are usually connected with various types of neighbors via different types of relations. As such, representation learning on heterogeneous graphs is much more challenging. It needs to not only incorporate heterogeneous structure (graph) information but also consider the heterogeneous attributes associated with each node [21].
Zhang et al. proposed a heterogeneous GNN model (HetGNN) for capturing both structure and content heterogeneity. HetGNN first captures the strongly correlated heterogeneous neighbors of each node and then aggregates feature information of these sampled neighbors [22]. Zhao et al. proposed a GNN-based framework named HGSL that jointly performs heterogeneous graph structure learning and GNN parameter learning for classification. In HGSL, the feature similarity graphs, the feature propagation graphs, and the semantic graphs are generated separately so as to comprehensively learn an optimal heterogeneous graph [23]. Zhang et al. designed a relation-centered pooling and convolution (PC-HGN) operation that enables relation-specific sampling and cross-relation convolutions on heterogeneous graphs, from which the structural heterogeneity of the graph can be better encoded into the embedding space through the adaptive training process [7]. HAN (Heterogeneous graph Attention Network) [24] is a heterogeneous graph learning framework based on node-level and semantic-level attention mechanisms. Specifically, node-level attention learns the importance between a node and its meta-path based neighbors, while semantic-level attention learns the importance of different meta-paths. Hu et al. presented a heterogeneous graph transformer (HGT) for modeling heterogeneous graphs [25]. HGT introduces the node-type and edge-type dependent attention mechanism while different trainable parameters are assigned to each node and edge type. HGT can incorporate high-order heterogeneous neighbor information, which automatically learns the importance of implicit meta path. Considering relation-aware characteristics, Yu et al. proposed a relation-aware representation learning model for heterogeneous graphs (R-HGNN). This model derives a fine-grained representation from a group of relation-specific node representations reflecting the characteristics of the node associated with a specified relation [26].
II-C Hypergraph Learning
Hypergraph learning explores the high-order correlations in the data, which extends the traditional graph learning models to a high dimensional and more complete nonlinear space. It offers a promising solution for analyzing complex structured data with satisfactory performance in practice [16]. The construction of the hypergraph from the given data is a key step for hypergraph learning, which significantly affects the final learning performance.
Zhang et al. proposed dynamic hypergraph structure learning (DHSL) to update hypergraph structure iteratively during the learning process [27]. It is essential to make dynamic modifications to the initial hypergraph structures from adjusted feature embedding. Feng et al. presented a framework named hypergraph neural network (HGNN) for handling complex and high-order correlations. In HGNN, the complex data correlation is formulated in a hypergraph structure, and a hyperedge convolution operation is used to exploit the high-order data correlation for representation learning [9].
To exploit high-order relations among the features, Jiang et al. proposed a dynamic hypergraph neural networks (DHGNN) framework that is composed of dynamic hypergraph construction (DHG) and hypergraph convolution (HGC). DHG utilizes the k-NN method to generate the basic hyperedge and extends the adjacent hyperedge set by k-means clustering, with which the local and global relations can be extracted. HGC is designed to encode high-order data relations in the hypergraph structure [28]. Cai et al. introduced a hypergraph structure learning (HSL) framework to optimize the hypergraph structure and HGNNs simultaneously in an end-to-end way. To efficiently learn the hypergraph structure, HSL adopts a hyperedge sampling strategy to prune the redundant hyperedges, which is followed by an incident node sampling for pruning irrelevant incident nodes and discovering potential implicit connections. The consistency between the optimized structure and the original structure is maintained by the intra-hyperedge contrastive learning module [29]. Gao et al. proposed a tensor-based dynamic hypergraph learning (t-DHL) model to efficiently learn dynamic hypergraphs. t-DHL utilizes a tensor representation to characterize the dynamic hypergraph structure more flexibly. During the optimization of the tensor representation, not only the weights but also the number and order of hyperedges can be adjusted. As an extended version of HGNN, a general hypergraph neural network framework named HGNN+ was proposed for modeling high-order representation among the data, which is achieved by bridging multi-modal/multi-type data and hyperedge with hyperedge groups [30].
Recent works mostly focus on neighborhood-based construction strategies when constructing hyperedges in the hypergraph. Specifically, hyperedge is often constructed through clustering a master vertex and its neighbor vertices. However, for the commonly adopted clustering methods like k-NN [12, 28, 9], they treat hypergraph construction and graph learning separately. As a result, these models are susceptible to noisy data which limits their applications. Apart from that, most neighborhood-based models fail to incorporate the heterogeneity of the topology, leading to a single type hyperedge [16]. In this work, we propose a heterogeneous hypergraph learning framework LFH to solve the above mentioned issues. LFH constructs the hyperedge based on different edge types, which can then be integrated into the learning process to improve the learned representations.
III Preliminaries and problem formulation
In this section, we first introduce the relevant definitions and then formulate the key problems associated with representation learning on heterogeneous hypergraphs. The notations used in this paper are listed in Table I.
Definition 1.
(Heterogeneous Pairwise Graph). A heterogeneous graph $\dot{\mathcal{G}}=\{\dot{\mathcal{V}},\dot{\mathcal{E}},\dot{\mathcal{T}_{e}},\dot{\mathcal{T}_{v}}\}$ is defined as a graph with multiple node types $\dot{\mathcal{T}_{v}}$ or edge types $\dot{\mathcal{T}_{e}}$, where $\dot{\mathcal{V}}=\{\emph{v}_{1},\emph{v}_{2},...,\emph{v}_{N}\}$ denotes the set of nodes and ˙=$\dot{\mathcal{E}}=\left\{\left(\emph{v}_{i},\emph{v}_{j}\right)|\emph{v}_{i},\emph{v}_{j}\in\dot{\mathcal{V}}\right\}$ denotes the set of edges. We define the node type mapping function $\dot{\phi}:\dot{\mathcal{V}}\rightarrow\dot{\mathcal{T}_{v}}$ and the edge type mapping function $\dot{\psi}:\dot{\mathcal{E}}\rightarrow\dot{\mathcal{T}_{e}}$, respectively.
Definition 2.
Figure 2: An illustrative example of the ACM dataset: (a) Heterogeneous pairwise graph including four node types (Author, Paper, Subject) and four edge types (Write, Written by, Belong to, Contain). (b) A heterogeneous hypergraph that models two implicit data relations: preference for subjects and the relation among the subjects.
[…]
$d\left(\emph{v}_{i}\right)=\sum_{\emph{e}_{p}\in\mathcal{E}}w\left(\emph{e}_{p}\right)\cdot\emph{H}\left(\emph{v}_{i},\emph{e}_{p}\right)$ (1)
$\delta\left(\emph{e}_{j}\right)=\sum_{\emph{v}_{p}\in\mathcal{V}}\emph{H}\left(\emph{v}_{p},\emph{e}_{j}\right)$ (2)
Figure 2 provides an illustrative comparison between a heterogeneous pairwise graph and a hypergraph. Figure 2 (a) sketches the ACM dataset composed of three node types (Author, Paper, Subject) and four edge types (Write, Written by, Belong to, Contain). By contrast, a heterogeneous hypergraph is capable of modeling a more complex graph structure by constructing different types of hyperedges, in which every hyperedge can connect any number of nodes to model a specific type of data relationship. As shown in Figure 2 (b), the purple hyperedge models an author’s preference for subjects, while the yellow one models the relation among different subjects.
Figure 3: An overview of our proposed LFH framework.
Problem 1.
(Hyperedge Generation in Heterogeneous Hypergraph). While node set $\mathcal{V}$ explicitly exists in graph data, the hyperedge set $\mathcal{E}$ is normally implicit. Following Definition 2, any hyperedge $\emph{e}\in\mathcal{E}$ contains a subset of nodes $\mathcal{V}$. We aim to learn a hyperedge construction function $f_{con}:\mathcal{V}\rightarrow\mathcal{E}$, where hyperedges are constructed based on the given nodes. Moreover, this function can be used for constructing incidence matrix H.
Problem 2.
(Representation Learning on Heterogeneous Hypergraph). Given a heterogeneous hypergraph $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathcal{T}_{e},\mathcal{T}_{v},\mathcal{W}\}$, we aim to learn a mapping function $f_{emb}\left(\mathcal{G}\right)\rightarrow\emph{Z}^{\mathcal{G}}$, where $\emph{Z}^{\mathcal{V}}\in\mathcal{R}^{d\times N}$ indicates representation embedding of all nodes in $\mathcal{G}$. This representation embedding can be used for downstream predictive tasks such as node classification and link prediction.
TABLE I: Table of key notations
Notation  Description
[…]
$\|\cdot\|_{1}$ $l_{1}$ norm
$\|\cdot\|_{2}$ $l_{2}$ norm
||$\cdot$|| ⋅ | The set size
$\mathcal{L}_{recon}$ Reconstruction loss
$\mathcal{L}_{label}$ Label loss
$Softmax\left(\cdot\right)$ Softmax function
$f_{con}$ Hyperedge construction function
$f_{emb}$ Representation embedding mapping function
$\Theta_{p}$,$\theta$,$\Theta^{att}$ Learnable model parameters
$\dot{\mathcal{G}}$ Heterogeneous pairwise graph
$\dot{\mathcal{V}}$ The set of nodes in the heterogeneous pairwise graph
$\dot{\mathcal{E}}$ The set of edgees in the heterogeneous pairwise graph
$\dot{\mathcal{T}_{v}}$ The set of node types in the heterogeneous pairwise graph
$\dot{\mathcal{T}_{e}}$ The set of edge types in the heterogeneous pairwise graph
$\dot{\emph{X}}$ Raw node features in the heterogeneous pairwise graph
$\dot{\emph{x}_{i}}$ Raw feature of node $v_{i}$ in the heterogeneous pairwise graph
$\dot{\phi}$ Node type mapping function in the heterogeneous pairwise graph
$\dot{\psi}$ Edge type mapping function in the heterogeneous pairwise graph
$\mathcal{G}$ Heterogeneous hypergraph
$\mathcal{V}$ The set of nodes in the heterogeneous hypergraph
$\mathcal{E}$ The set of edgees in the heterogeneous hypergraph
$\mathcal{T}_{v}$ The set of node types in the heterogeneous hypergraph
$\mathcal{T}_{e}$ The set of edge types in the heterogeneous hypergraph
$\phi$ Node type mapping function in the heterogeneous hypergraph
$\psi$ Edge type mapping function in the heterogeneous hypergraph
X Initial node embedding in the heterogeneous hypergraph
$\emph{x}_{i}$ Initial node embedding of node $v_{i}$ in the heterogeneous hypergraph
$\mathcal{W}$ Weight matrix of hyperedges
H Incidence matrix
$\emph{Z}^{\mathcal{V}}$ Embedding of all nodes in $\mathcal{V}$
$\emph{Y}^{\mathcal{V}}$ Labels of nodes
$v$ A node $v\in\mathcal{V}$
$e$ A node $e\in\mathcal{E}$
$w\left(e\right)$ The weight of hyperedge $e$
$d\left(v\right)$ Degree of node $v$
$\delta\left(e\right)$ Degree of edge $e$
$v_{i}^{m}$ The master node of $v_{i}^{m}$
$\emph{S}\left(v_{i}^{m}\right)$ The slave node set of the master node $v^{m}_{i}$
$\tilde{\emph{S}}\left(v_{i}^{m}\right)$ The candidate slave node set of the master node $v^{m}_{i}$
$\mathcal{T}\left(\tilde{\emph{S}}\left(v_{i}^{m}\right)\right)$ The type of the candidate slave node set
$\hat{\emph{X}}\left(v_{i}^{m}\right)$ The reconstructed vector of the master node $v_{i}^{m}$
$\emph{p}_{v_{i}^{m}}$ The reconstruction coefficient vector of $v_{i}^{m}$
$c\left(v^{m}_{i}\right)$ Reconstruction loss of the master node $v^{m}_{i}$
$M$ The number of hyperedges
$N$ The number of nodes
$n_{i}$ The number of candidate slave node set associated with node $v_{i}$
$M_{i}$ The number of hyperedges associated with the node $v_{i}$
E Hyperedges embeddding
$K$ The number of attention heads
IV Heterogeneous Hypergraph
In this section, we introduce our proposed Learning from Heterogeneity (LFH) framework for hypergraphs. Figure 3 provides an overview of LFH. We first demonstrate the pairwise fusion process from the heterogeneous pairwise graph to obtain a high-quality initial node embedding. Afterward, we dynamically generate the heterogeneous hyperedges among the nodes with different node types through feature reconstruction. After the heterogeneous hypergraph is constructed, our proposed dynamic hypergraph learning can then be performed to train the node embedding space, in which node representations can be derived for downstream node classification and link prediction tasks.
IV-A Initial Node Embedding
As demonstrated in Fig. 3, the input of this step is the heterogeneous pairwise graph. It is of great importance to generate high-quality initial node embedding to better exploit the implicit data relations in the hypergraph. Intuitively, fusing pairwise information from a heterogeneous pairwise graph into the initial node embeddings can help capture high-order correlations among the nodes and thus affects the construction of the hypergraph. Given a heterogeneous pairwise graph $\dot{\mathcal{G}}=\{\dot{\mathcal{V}},\dot{\mathcal{E}},\dot{\mathcal{T}_{e}},\dot{\mathcal{T}_{v}}\}$ and its raw node features $\dot{\emph{X}}=\left[\dot{\emph{x}}_{1},\dot{\emph{x}}_{2},...,\dot{\emph{x}}_{N}\right]\in\mathcal{R}^{D\times N}$, where the raw feature of node $v_{i}\in\mathcal{V}$ is $\dot{\emph{x}}_{i}\in\mathcal{R}^{D}$. The pairwise fusion can be implemented by any GNN model that operates on heterogeneous pairwise graphs. Formally, the pairwise fusion process can be defined as:
[…]
IV-B Heterogeneous Hyperedge Generation
The special structure of hyperedges enables hypergraphs to encode high-order data relations. By encircling a specific set of nodes with common attributes or close relationships within one hyperedge, we can effectively represent local group information among the nodes. Therefore, it is essential to discover the related nodes for different hyperedges. To facilitate hyperedge construction, we define two types of nodes: master node and slave node. A master node $v^{m}_{i}\in\mathcal{V}$ serves as an anchor when constructing a hyperedge, which dynamically encircles a set of slave nodes $\{v^{s}_{i}\}$ to jointly form the hyperedge. The collection of participant nodes is denoted as $S\left(v^{m}_{i}\right)$, which is united with master node $v_{i}^{m}$ to represent a complete data relation within a hyperedge. Previous works have adopted the $k$-NN method to select the k-th nearest nodes to form a hyperedge [27, 31]. Alternative approaches solve an independent optimization problem to generate the hyperedges [12, 32], however, this solution fails to incorporate the heterogeneity of the hyperedges, and thus cannot be integrated into the representation learning process.
To address the above challenges, here we propose a dynamic learning process that generates heterogeneous hyperedges and adaptively reconstructs the hyperedge in different types. Specifically, for each master node $v^{m}_{i}\in\mathcal{V}$, our framework encloses the related slave nodes of various types to encode heterogeneous topological structures with implicit relationships, based on different attributes among the nodes. Suppose the master node $v^{m}_{i}$ has a set of $n_{i}$ candidate slave nodes $\tilde{S}^{k}\left(v^{m}_{i}\right)$, where $k=\left[1,...,n_{i}\right]$ with different types. The hyperedge can be reconstructed based on the node types of the master node and candidate slave node sets. The type of the candidate slave node set can be defined as $\mathcal{T}\left(\tilde{S}^{k}\left(v^{m}_{i}\right)\right)$ and clearly is a subset of node type $\mathcal{T}_{v}$. Each candidate slave node set $\tilde{S}^{k}\left(v^{m}_{i}\right)$ contains all nodes with one or several node types, which is denoted as $\tilde{S}^{k}\left(v^{m}_{i}\right)=\left\{v|\phi\left(v\right)\in\mathcal{T}\left(\tilde{S}^{k}\left(v^{m}_{i}\right)\right),v\neq v_{i}^{m}\right\}$. Since each hyperedge is generated from the corresponding candidate slave node set of a certain type, the type number of the hyperedge is equal to the type number of the candidate slave node set. An exemplary hyperedge construction is demonstrated in the middle part of Figure 3, where the master node encompasses $|\mathcal{T}_{v}|$ candidate slave node sets with distinct types. For each type, a hyperedge (represented as the coloured triangle) is dynamically constructed. Mathematically, this process of generating a hyperedge from a master node $v_{i}^{m}$ and one of its candidate slave node set $\tilde{S}^{k}\left({v_{i}^{m}}\right)$ can be denoted as:
$\hat{\emph{X}}_{k}\left(v^{m}_{i}\right)=\emph{X}\left(\tilde{S}^{k}\left(v^{m}_{i}\right)\right)\cdot\emph{p}_{v^{m}_{i}}^{k}$ (4)
It can be interpreted as the selection of slave nodes for this hypergraph through a linear combination of candidate slave node embedding $\emph{X}\left(\tilde{S}^{k}\left(v^{m}_{i}\right)\right)\in\mathcal{R}^{d\times|\tilde{S}^{k}\left(v^{m}_{i}\right)|}$ and trainable reconstruction coefficient vector $\emph{p}_{v^{m}_{i}}^{k}\in\mathcal{R}^{|\tilde{S}^{k}\left(v^{m}_{i}\right)|}$. Each element $p_{v^{m}_{i}}^{k}\left(v\right)$ is the learned reconstruction coefficient associated with the node $v\in\tilde{S}^{k}\left(v^{m}_{i}\right)$. Based on $\emph{p}_{v^{m}_{i}}^{k}$, the nodes in the candidate slave node set $\tilde{S}^{k}\left(v^{m}_{i}\right)$ with reconstruction coefficient larger than zero are grouped to generate a hyperedge with the master node, which is denoted as $S^{k}\left(v^{m}_{i}\right)=\left\{v|v\in\tilde{S}^{k}\left(v^{m}_{i}\right),\emph{p}_{v^{m}_{i}}^{k}\left(v\right)>0\right\}$. The reconstruction error $c^{k}\left(v^{m}_{i}\right)$, which measures the difference between the master node and the reconstructed master node, is calculated as follows:
$c^{k}\left(v^{m}_{i}\right)=\|\hat{\emph{X}}_{k}\left(v^{m}_{i}\right)-\theta_{k}\cdot\emph{X}\left(v^{m}_{i}\right)\|_{2}$ (5)
$\theta_{k}\in\mathcal{R}^{d\times d}$ is a type-specific trainable projection matrix and $\|\cdot\|_{2}$ is the $l_{2}\text{-}norm$ regularization of a vector. Note that the purpose of hyperedge construction is to dynamically group the most relevant nodes, forming a hyperedge in which the nodes have strong dependencies. Practically, we add $l_{1}\text{-}norm$ regularization and $l_{2}\text{-}norm$ regularization of p simultaneously. The $l_{1}\text{-}norm$ regularization constrains the reconstruction coefficient to zero, which tends to group fewer slave nodes to reconstruct the master node. However, the $l_{1}\text{-}norm$ regularization is sensitive to noise and outliers, which leads to non-smooth results. For this reason, we consider the $l_{1}\text{-}norm$ regularization and $l_{2}\text{-}norm$ regularization simultaneously and use a weight hyperparameter $\gamma$ to trade off the effect of both. Overall, the loss of this process is defined as:
$\mathcal{L}_{recon}=\sum_{i=\left[1,...,N\right]}\sum_{k=\left[1,...,n_{i}\right]}\lambda c^{k}\left(v^{m}_{i}\right)+\|\emph{p}_{v^{m}_{i}}^{k}\|_{1}+\gamma\|\emph{p}_{v^{m}_{i}}^{k}\|_{2}~{}$ (6)
where $\|\cdot\|_{1}$ denotes the $l_{1}\text{-}norm$ regularization of a vector, $\lambda$ is the weight hyperparameter of the reconstruction error, and $\gamma$ is the norm hyperparameter to trade off the $l_{1}\text{-}norm$ regularization and $l_{2}\text{-}norm$ regularization of the reconstruction coefficient vector $\emph{p}_{v^{m}_{i}}^{k}$.
From the $n_{i}$ candidate slave node sets of the master node $v_{i}^{m}$, we thus can generate the $n_{i}$ hyperedges of different types. Iteratively, each node $v_{j}$ could act as the master node to generate several hyperedges of different types to lead to the final hypergraph. The corresponding incidence matrix H of the hypergraph is denoted as:
$\emph{H}\left(v_{j},e^{k}\left(v_{i}^{m}\right)\right)$ $=$ $\begin{cases}1&\text{$v_{j}=v_{i}^{m}$}\\ \emph{p}_{v^{m}_{i}}^{k}\left(v_{j}\right)&\text{$v_{j}\in e^{k}\left(v^{m}_{i}\right),v_{j}\neq v_{i}^{m}$}\\ 0&\text{otherwise}\end{cases}$ (7)
Figure 4: Type-specific multi-head attention on node embedding update
IV-C Dynamic Hypergraph Learning
As illustrated in Fig 3, dynamic hypergraph learning consists of two key components: hyperedge embedding updating and multi-head attention node embedding updating. This section elaborates on these components and loss function.
IV-C1 Hyperedge Embedding Updating
We aggregate node embeddings to the hyperedge containing these nodes. The hyperedges embedding is denoted as $\emph{E}\in\mathcal{R}^{d\times M}$. The embedding of each hyperedge $e^{k}\left(v^{m}_{i}\right)$ is obtained as follows:
$\emph{E}\left(e^{k}\left(v^{m}_{i}\right)\right)=\frac{\emph{X}\cdot\emph{H}\left(e^{k}\left(v^{m}_{i}\right)\right)}{\delta\left(e^{k}\left(v^{m}_{i}\right)\right)}$ (8)
$\emph{H}\left(e^{k}\left(v^{m}_{i}\right)\right)\in\mathcal{R}^{N\times 1}$ is a column of incidence matrix H (See Eq. 7), where each element is a coefficient that represents the importance of the node to the hyperedge $e^{k}\left(v^{m}_{i}\right)$. The degree of edge $\delta\left(e^{k}\left(v^{m}_{i}\right)\right)$ is the normalization factor in the process of hyperedge embedding updating.
IV-C2 Multi-head Attention Node Embedding Updating
Previous works calculate the similarities among hyperedges and then use them as weighting coefficients when updating respective node embeddings [33, 32]. However, the heterogeneity in the hypergraph is omitted, especially with multiple types of hyperedge and nodes. To cope with that, we design a heterogeneous multi-head attention mechanism that can train the importance of heterogeneous hyperedges with respect to nodes as shown in Figure 4.
We use all hyperedges associated with master node $v^{m}_{i}$ to update the node embedding. The hyperedges associated with node $v^{m}_{i}$ is represented as $\mathcal{E}\left(v^{m}_{i}\right)=\left\{e^{1}\left(v^{m}_{i}\right),e^{2}\left(v^{m}_{i}\right),...,e^{M_{i}}\left(v^{m}_{i}\right)\right\}$.
Figure 4 illustrates the calculation process of a master node $v^{m}_{i}$ and two related hyperedges $e^{1}\left(v^{m}_{i}\right)$ and $e^{2}\left(v^{m}_{i}\right)$: we first calculate the multi-head attention between $v_{i}^{m}$ and $e^{1}\left(v^{m}_{i}\right)$, and then use the normalized attention as the weight of $e^{1}\left(v^{m}_{i}\right)$. The calculation involves Eq.9 to Eq.13.
$\emph{Q}^{h}\left(v^{m}_{i}\right)=Q\text{-}Proj^{h}_{\phi\left(v^{m}_{i}\right)}\cdot\emph{X}\left(v^{m}_{i}\right)$ (9)
where the dimension of $\emph{X}\left(v^{m}_{i}\right)$ is $\mathcal{R}^{d\times 1}$, and $Q\text{-}Proj^{h}_{\phi\left(v^{m}_{i}\right)}\in\mathcal{R}^{\frac{d}{K}\times d}$ represents the projection matrix on the $h$-th attention head, $h\in\left[1,K\right]$. $K$ is the number of attention heads. Note the projection matrices are distinguished on hyperedge types. Similarly, we also project the hyperedge $e^{1}\left(v^{m}_{i}\right)$ through $K\text{-}Proj^{h}_{\psi\left(e^{1}\left(v^{m}_{i}\right)\right)}$ into $\emph{K}^{h}\left(e^{1}\left(v^{m}_{i}\right)\right)$ as:
$\emph{K}^{h}\left(e^{1}\left(v^{m}_{i}\right)\right)=K\text{-}Proj^{h}_{\psi\left(e^{1}\left(v^{m}_{i}\right)\right)}\cdot\emph{E}\left(e^{1}\left(v^{m}_{i}\right)\right)$ (10)
where $\emph{E}\left(e^{1}\left(v^{m}_{i}\right)\right)\in\mathcal{R}^{d\times 1}$ is the embedding of hyperedge $e^{1}\left(v^{m}_{i}\right)$. The dimension of $K\text{-}Proj^{h}_{\psi\left(e^{1}\left(v^{m}_{i}\right)\right)}$ is $\mathcal{R}^{\frac{d}{K}\times d}$ while the dimension of the output $\emph{K}^{h}\left(e^{1}\left(v^{m}_{i}\right)\right)$ is $\mathcal{R}^{\frac{d}{K}}$.
$\begin{split}att^{h}(v^{m}_{i},e^{1}(v^{m}_{i}))=\big{(}\emph{Q}^{h}&(v^{m}_{i})\cdot\Theta^{att}_{\psi\left(e^{1}\left(v^{m}_{i}\right)\right)}\\ &\cdot\emph{K}^{h}\left(e^{1}\left(v^{m}_{i}\right)\right)\big{)}\cdot\frac{\mu_{\psi(e^{1}(v^{m}_{i}))}}{\sqrt{d}}\end{split}$ (11)
Eq. 11 derives the attention value of the $h$-th head between $v^{m}_{i}$ and $e^{1}\left(v^{m}_{i}\right)$. The type-specific learnable matrix $\Theta^{att}_{\psi\left(e^{1}\left(v^{m}_{i}\right)\right)}\in\mathcal{R}^{\frac{d}{K}\times\frac{d}{K}}$ for edge type $\psi\left(e^{1}\left(v^{m}_{i}\right)\right)$ represents the learnable semantic information for each edge type. $\mu$ is a scaling factor for different hyperedge types. Moreover, since the magnitude of K and Q can increase the attention value significantly and eventually lead to the gradient explosion problem, we divide the obtained value by $\sqrt{d}$ to stabilize the training process.
Following that, the calculated attention value $att^{h}\left(v^{m}_{i},e^{1}\left(v^{m}_{i}\right)\right)$ is able to quantify the importance of $e^{1}\left(v^{m}_{i}\right)$ to $v^{m}_{i}$ on the $h$-th adaptively. The weight of hyperedge $e^{1}\left(v^{m}_{i}\right)$ is obtained as follows:
$\tilde{\emph{w}}\left(v^{m}_{i},e^{1}\left(v^{m}_{i}\right)\right)=\mathop{\operatorname*{\mathchoice{\Big{\|}}{\big{\|}}{\|}{\|}}}\limits_{h\in\left[1,K\right]}att^{h}\left(v^{m}_{i},e^{1}\left(v^{m}_{i}\right)\right)$ (12)
[…]
$\emph{w}\left(v^{m}_{i},e\right)=\mathop{Softmax}\limits_{\forall e\in\mathcal{E}\left(v^{m}_{i}\right)}\left(\tilde{\emph{w}}\left(v^{m}_{i},e\right)\right)$ (13)
where $\emph{w}\left(v^{m}_{i},e\right)=\left[\emph{w}\left(v^{m}_{i},e^{1}\left(v^{m}_{i}\right)\right),...,\emph{w}\left(v^{m}_{i},e^{M_{i}}\left(v^{m}_{i}\right)\right)\right]\in\mathcal{R}^{K\times M_{i}}$, with entries are normalized among all hyperedges per head (illustrated in the right part of Fig. 4). The attentive aggregation of head $h$ for node $v^{m}_{i}$ is formulated as:
$\emph{z}_{i}^{h}=\sum_{k\in\left[1,M_{i}\right]}\left(w^{h}\left(v^{m}_{i},e^{k}\left(v^{m}_{i}\right)\right)\times\emph{K}^{h}\left(e^{k}\left(v^{m}_{i}\right)\right)\right)$ (14)
$w^{h}\left(v^{m}_{i},e^{k}\left(v^{m}_{i}\right)\right)$ is the scalar element of $h$-th head in the weight vector $\emph{w}\left(v^{m}_{i},e^{k}\left(v^{m}_{i}\right)\right)$ of hyeredge $e^{k}\left(v^{m}_{i}\right)$, and the dimension of $\emph{z}_{i}^{h}$ is $\mathcal{R}^{\frac{d}{K}}$. After concatenating all $K$ heads, we obtain the updated node embedding of $v^{m}_{i}$ through a shallow multi-layer perceptron (MLP), which is denoted as:
[…]
IV-D Loss Function
We realize LFH in an end-to-end fashion, including the initialization of node embedding, the dynamic heterogeneous hyperedges generation and heterogeneous hyperedges learning. The learnt node representations can then be used in downstream tasks such as node classification and link prediction. The integrated training process is conducted in a supervised manner. We propose a united loss function, which combines the loss of reconstruction in the dynamic heterogeneous hyperedges generation and the loss of supervised downstream task. We take node classification as an example of a downstream task. In this case, the label of the node set $\mathcal{V}$ is $\emph{Y}^{\mathcal{V}}$. We use an MLP to map node representations to label space, and the loss in node classification can be calculated as:
[…]
where CE is the cross-entropy function and used to measure the differences between estimated labels and the true labels of nodes. Finally, the united loss is defined as:
$\mathcal{L}_{u}=\left(1-\alpha\right)\times\mathcal{L}_{label}+\alpha\mathcal{L}_{recon}$ (17)
where $\mathcal{L}_{recon}$ is defined in Eq. 6. The process of reconstructing the master node using slave nodes can uncover implicit relationships between the master node and others. However, as pointed out above, this process is sensitive to noisy nodes. Over-reliance on reconstruction loss compromises the performance of our model. To enhance the generalization ability, we need to use a weight hyperparameter $\alpha$, to balance the effects of loss of supervised downstream task and reconstruction loss.
IV-E Model Analysis and Discussion
We analyze the computational cost and summarize the advantages of LFH as below:
IV-E1 Computational Cost
In the process of heterogeneous hypergraph generation, the computational cost mainly lies in the reconstruction of the master node from its candidate slave node set (Eq. 4). As it is essential to generate hyperedges of different types for each node, and the computational cost of heterogeneous hypergraph generation is $\emph{O}\left(d|\mathcal{T}_{v}|N\right)$, where $N$ is the number of nodes, $d$ is the dimensionality of node embedding, and $|\mathcal{T}_{v}|$ is the number of node types. Additionally, for the process of dynamic hypergraph learning, we use the incidence matrix H and nodes embedding X to update hyperedges embedding, with the computational cost of $\emph{O}\left(dMN\right)$. Note the number of hyperedges $M$ is proportional to the product of the $N$ and the number of edge types $|\mathcal{T}_{e}|$, thus the computational cost of hyperedges embedding updating is $\emph{O}\left(d|\mathcal{T}_{e}|N^{2}\right)$. Moreover, the computational cost of multi-head attention node embedding updating is $\emph{O}\left(d^{2}|\mathcal{T}_{e}|N\right)$, where $\mathcal{T}_{e}$ is the number of hyperedge types. Hence, the total computational cost of our proposal is $\emph{O}\left(d|\mathcal{T}_{e}|N^{2}+\left(d|\mathcal{T}_{v}|+d^{2}|\mathcal{T}_{e}|\right)N\right)$. The scale of $d$, $\mathcal{T}_{e}$ and $\mathcal{T}_{v}$ are usually far less than $N$, This computational cost is on par with many existing hypergraph-based models, such as [12] and [34].
IV-E2 Discussion
Compared to pairwise graph learning models such as GAT [6] and PC-HGN [7], LFH is capable of modeling implicit high-order data relations. Moreover, LFH has two advantages. First, the process of hyperedges generation is conducted with hypergraph learning within a united training process while taking into account the heterogeneity of the graph. This enables an adaptive hypergraph modeling than HGNN [9], whose hyperedges are generated by clustering method k-NN. Other works that use static clustering methods for hyperedge generation adopt different strategies of k-means or a combination of k-NN and k-means to generate different hyperedges [27, 35]. The main limitation of this line of work is the static clustering methods are sensitive to noise and outliers. Another issue is the hyperparameters of clustering method may affect the hyperedge generation, and adaptively calibrating the hyperparamters of clustering method during training is non-trivial. LFH, on the other hand, can incorporate hyperedges information of different types in different heads, using a heterogeneous multi-head attention mechanism. The mechanism dynamically quantifies the weights of hyperedges based on the correlation between the node representation and the related hyperedges representation, making the weights more descriptive of the importance of the related hyperedges to the node.
V Experiment
In this section, we empirically evaluate the effectiveness of our LFH framework and analyze the impacts of its key components to the final performance. We conduct our experiments on three different graph datasets, including DBLP, IMDB, and ACM. Additionally, we compare the performance of LFH with eleven baseline models, which include homogenous graph learning, heterogeneous graph learning, and hypergraph learning models.
V-A Experimental Setup
TABLE II: Statistics of the datasets.
Dataset # Nodes # Node Types # Edges # Edge Types # Node Features
ACM 8994 3 12961 4 1902
DBLP 18405 3 33973 4 334
IMDB 12772 3 18644 4 1256
V-A1 Dataset
The experiments are conducted on three datasets, which are popularly used in classification tasks. The details of these datasets are summarized in Table II.
  • DBLP [^1] https://s3.cn-north-1.amazonaws.com.cn/dgl-data/dataset/openhgnn/dblp4GTN.zip is an academic network from four research areas, including database, machine learning, data mining and information retrieval. It uses Paper (P), Author (A), and Conference (C) as different node types, while edges are presented as P-A, A-P, P-C and C-P with different edge types. Four research areas are used as labels for this dataset. The initial node features are calculated using bag-of-words.
  • ACM222 https://s3.cn-north-1.amazonaws.com.cn/dgl-data/dataset/openhgnn/acm4GTN.zip shares a similar data characterises with DBLP. It contains Paper (P), Author (A) and Subject (S) as node types, along with four types of edges (P-A, A-P, P-S and S-P). The papers are labelled into three classes (Database, Wireless Communication, Data Mining). It also uses bag-of-words to construct the initial node features.
  • IMDB333 https://s3.cn-north-1.amazonaws.com.cn/dgl-data/dataset/openhgnn/imdb4GTN.zip contains movie (M), actors (A) and directors (D). Each movie is labelled according to its genre (Action, Comedy, Drama). Node features are also initialized using bag-of-words.
V-A2 Baseline Models and Configurations
We compare with some state-of-the-art baselines, each of which reports promising results in the classification task. Although these baselines share the same goal of learning the representation of the graph, they are originally designed to fulfil graph learning on different graph data, including homogeneous pairwise graph learning, pairwise heterogeneous graph learning, and hypergraph learning.
  • GCN [5] is a graph convolutional network designed specifically for homogeneous pairwise graph learning. The depth of the layer in GCN is set to 2.
  • GAT [6] is the first work that introduces the attention mechanism in homogeneous graph learning. It enables the weighted message aggregations from neighbour nodes. The number of attention heads is set to 3.
  • GraphSAGE [17] designs a sampling approach when aggregating messages from neighbour nodes. It also supports different aggregation functions. The sample window of GraphSAGE is set to 10.
  • GraphSAINT [20] splits nodes and edges from a bigger graph into a number of subgraphs, on which the GCN is applied for node representation learning. We adopt the node sampling strategy for GraphSAINT and use 8000 as the node budget and 25 as the number of subgraphs.
  • HAN [24] uses attention techniques on heterogeneous graph learning, in which the node embeddings are updated through manually designed meta-paths. The number of attention heads is set to 8.
  • HGT [25] designs type-specific attention layers, assigning different trainable parameters to each node and edge type. For the best performance of HGT, the number of attention heads is set to 8.
  • R-HGNN [26] proposes a relation-aware model to learn the semantic representation of edges while discerning the node representations with respect to different relation types. It is designed to learn heterogeneous graph representations. The depth of the layer in R-HGNN is set to 2.
  • PC-HGN [7] employs a sampling-based convolution. It designs an efficient cross-relation convolution that allows message aggregations of a node from different connected relation types simultaneously. We set the number of kernels to 64 and pooling size to 2 as the best performance reported.
  • HL [36] uses the hypergraphs to represent complex relationships among the objects, where where attribute values are regarded as hyperedges. The value of the regularization factor is set to 0.1.
  • HGNN [9] performs convolution with a hypergraph Laplacian, which is further approximated by truncated Chebyshev polynomials to handle the data correlation during representation learning. The number of neighbors for generating hyperedges is set to 10.
  • HGNN+ [30] utilizes the multi-modal/multi-type correlation from each modality/type, and conducts hypergraph convolution in the spatial domain to learn a general data representation for various task. The value of $k$ for selecting $k$-hop neighbors to generate hyperedges is set to 1.
The shared parameter setup of the framework can be found in Table III. We randomly split all datasets into train/validate/test with the ratio of 0.2/0.1/0.7, respectively. It is worth mentioning that the hyperparameters of all baseline models are set to the optimal values reported in their original papers. All models are trained with a fixed 100 epochs, using an early stopping strategy when the performance on the validation set is not improved for 30 consecutive epochs. All trainable parameters of the neural network are initialized through Xavier [37] and optimized using Adam [38] with the learning rate 2e-3. The dropout rate is set to be 0.3 [39]. The optimal selection of hyperparameters of our model is further analyzed in the following section.
TABLE III: Shared parameter setup
Name  Setup
Optimizer Adam [38]
Learning rate 2e-3
Dropout rate 0.3
Weight initializer Xavier [37]
Epoch 100
MLP 1 linear layer, 1 output layer
TABLE IV: Performance of different models on node classification task.
Dataset Data split ratio Hypergraph-based Heterogeous-based Homogeous-based
LFH HGNN HGNN+ HL PC-HGN HAN R-HGNN HGT GCN GAT GraphSAGE GraphSAINT
ACM 10% 0.818 0.557 0.655 0.471 0.768 0.551 0.512 0.517 0.505 0.547 0.548 0.512
30% 0.871 0.801 0.851 0.544 0.845 0.842 0.847 0.814 0.762 0.767 0.763 0.771
50% 0.882 0.804 0.852 0.596 0.849 0.842 0.854 0.828 0.846 0.853 0.851 0.839
70% 0.905 0.840 0.882 0.607 0.872 0.866 0.864 0.857 0.853 0.858 0.858 0.856
IMDB 10% 0.491 0.410 0.432 0.415 0.407 0.372 0.406 0.341 0.322 0.384 0.351 0.361
30% 0.539 0.442 0.463 0.424 0.459 0.398 0.415 0.388 0.384 0.429 0.388 0.409
50% 0.558 0.463 0.446 0.437 0.461 0.427 0.438 0.424 0.438 0.458 0.427 0.422
70% 0.571 0.494 0.492 0.459 0.488 0.464 0.487 0.435 0.445 0.488 0.454 0.460
DBLP 10% 0.861 0.669 0.819 0.503 0.818 0.812 0.672 0.598 0.535 0.574 0.543 0.597
30% 0.874 0.739 0.842 0.575 0.838 0.837 0.783 0.632 0.699 0.693 0.696 0.676
50% 0.888 0.767 0.866 0.634 0.861 0.855 0.853 0.724 0.755 0.775 0.771 0.776
70% 0.917 0.791 0.892 0.656 0.884 0.878 0.863 0.782 0.772 0.780 0.778 0.778
TABLE V: Performance of different models on link prediction.
Model ACM DBLP IMDB
Node2Vec [40] 0.689 0.628 0.724
GAT [6] 0.671 0.606 0.703
PC-HGN [7] 0.833 0.782 0.842
HGNN+ [30] 0.632 0.689 0.784
LFH 0.861 0.809 0.874
V-B Performance Analysis
We evaluate the performance of our proposed framework on node classification and link prediction tasks.
V-B1 Node Classification
In the node classification task, we split the original data into training data and test data. The training data are then split into a range of [10%, 30%, 50%, 70%]. The strategy of generating the training dataset is the same as [41], where each randomly selected node at least has one pairwise relation with others. The best F1 results among all datasets are demonstrated in Table IV, in which the highest scores are marked in bold. We compared three different categories of graph learning models: homogeneous pairwise graph learning models, heterogeneous pairwise graph learning models, and hypergraph learning models. Our proposed LFH achieves performance gain over other baseline models in all datasets by 2%-35.8%. Our model outperforms all baseline models in all datasets by an average improvement of 12.5%. Compared to four homogeneous pairwise GNN baselines GCN, GAT, GraphSAGE and GraphSAINT, the average F1 score improvement of LFH for all datasets is 14.4%. Regarding the cherry-picked heterogeneous pairwise graph learning models and hypergraph learning models, our model achieved an average improvement of 9.7% and 13.7%, respectively, compared to the best-recorded results in all datasets. It is worth noting that HGNN+ [30] is the only hypergraph learning baseline that utilizes pairwise information for hyperedge generation, resulting in better performance compared to other hypergraph learning models, except for our model. Furthermore, the training process of our model, which incorporates type-specific hyperedge generation and attentive embedding updates that exploits the heterogeneity of the graph, enables LFH to improve HGNN+ by an average of 5.7% in all datasets. Additionally, compared to the latest reporting heterogeneous learning model PC-HGN [7], the best results of our model gains an average increase by 5.2%. For different split ratios of training data, we observe that LFH still outperforms all baselines. When the split ratios are 70%, 50%, 30% and 10%, respectively, our model outperforms other models by 8.6%, 8.7%, 10.9% and 17.7% on average. The obtained result demonstrates that the robustness and generalization capability of our model in the informative case.
V-B2 Link Prediction
To investigate the effectiveness of our proposed framework, we also apply our model to the link prediction task. In this task, a graph with a certain faction of edges removed is given, and thus the objective is to predict these missing edges. Following the experiment setting adopted in [40], 50% percent of existing edges are randomly set hidden as positive samples while ensuring that the residual graph after the edge being removed remains connected. To generate negative examples, we randomly sample an equal number of node pairs from the graph, which have no edge connecting them. Similar to the related works [10, 42], we use a logistic regression classifier as a predictor for link prediction. To get the representation of each edge, we use Hadamard operator to compute the element-wise product for the linked node pairs. We compare the performance of LFH with four baselines. Node2vec [40] is a classic model for link prediction task. The other three baselines, GAT, PC-HGN, and HGNN+, are the top-performing baselines in homogenous pairwise graph learning models, heterogeneous pairwise graph learning models, and hypergraphs learning models in node classification tasks. Experimental results are summarized in Table V. LFH obtains performance gains of 16.8%, 18.8%, 2.9% and 14.6%, compared with Node2vec, GAT, PC-HGN, and HGNN+ and outperforms four baseline models in three datasets by an average improvement of 13.3%. It shows the promising capability of our model to capture implicit data relations. It is worth noting that HGNN+ that performs well in node classification does not have comparable performance in link prediction. This is because the HGNN+ applied for link prediction task heavily relies on pairwise edges to generate hyperedges, and thus a sharp decrease in the number of existing edges leads to a deterioration in performance.
V-C Impact of Hyperedge Construction
In our proposed LFH, $\lambda$ and $\gamma$ are used as hyperparameters. Specifically, $\lambda$ is the weight hyperparameter of the reconstruction error and $\gamma$ is the norm hyperparameter to trade off $l_{1}\text{-}norm$ regularization norm and $l_{2}\text{-}norm$ regularization of the reconstruction coefficient vector (See Eq. 6). Figure 5 and 6 illustrate the overall performance of LFH influenced by $\lambda$ and $\gamma$. Figure 5 illustrates the influence of $\lambda$ for different datasets on F1 score. In this figure, we fix $\gamma$ as 0.02, 0.2 and 0.2 while $\lambda$ ranges between 0.0001 and 1. When $\lambda$ is small, the reconstruction loss of $\lambda c^{k}\left(v^{m}_{i}\right)$ becomes trivial, and the reconstruction coefficient vector p will not experience large adjustments throughout the training phase, thus affecting the generation of hyperedges and deteriorating the performance of LFH. In contrast, when $\lambda$ increases to around 0.2, reconstruction loss of $\lambda c^{k}\left(v^{m}_{i}\right)$ accounts for the appropriate proportion of the total loss, thus achieving the best performance under this setup. When $\lambda$ continually increases to 20, F1 score drops rapidly, which clearly demonstrates that the dominance of reconstruction loss compromises the generalization ability of LFH.
On the other hand, Figure 6 describes the influence of $\gamma$. In this figure, we fix $\lambda$ as 0.001, 0.01 and 0.1, while $\gamma$ ranges between 0.002 and 20. When $\gamma$ is very small, The decisive role of $l_{1}\text{-}norm$ regularization enables our model to encircle fewer slave nodes to generate hyperedges, which makes our model more sensitive to noise and outliers. When $\gamma$ is large, the dominance of $l_{2}\text{-}norm$ regularization leads to the over-smoothing problem. With $\gamma$ ranging in $\left[0.0001,1\right]$, the F1 score reaches the highest when $\gamma=0.2$ for all fixed value of $\lambda$ in all datasets, and decreases afterwards. The result demonstrates that considering $l_{1}\text{-}norm$ regularization norm and $l_{2}\text{-}norm$ regularization simultaneously can improve the generalization ability of our model.
(a) ACM
(b) DBLP
(c) IMDB
Figure 5: Performance of LFH as a function of $\lambda$ for several values of $\gamma$ for different datasets.
(a) ACM
(b) DBLP
(c) IMDB
Figure 6: Performance of LFH as a function of $\gamma$ for several values of $\lambda$ for different datasets.
V-D Sensitivity Analysis
We study the impact of key parameters in LFH including the node embedding size and the number of attention head as illustrated in Table VI and VII.
V-D1 Impact of the node embedding size
We test the impact of different node embedding sizes for all datasets with data split ratio of 10%. As shown in Table VI, the performance of the model increases at first and then starts to drop, reporting the best result with the embedding size 256. Intuitively, larger dimension sizes of node embedding bring in extra redundancies in the training process, thus causing unexpected performance drops.
V-D2 Impact of the number of muti-head
We analyze the effect of the number of multi-head $K$. As observed in Table VII, when the number of attention heads increases, there is typically an improvement in the performance of our model. However, it is observed that the performance of LFH improves only marginally afterwards while causing large computational costs. When $K=4$, it is an optimal point to trade off the performance and computational cost.
TABLE VI: Performance of LFH with different node embedding sizes.
Initial Node Embedding Size ACM DBLP IMDB
128 0.800 0.818 0.456
256 0.818 0.861 0.491
512 0.789 0.766 0.424
TABLE VII: Performance of LFH with different number of attention heads.
Attention Head Number ACM DBLP IMDB
2 0.801 0.850 0.472
4 0.818 0.861 0.491
8 0.819 0.862 0.491
Figure 7: Ablation study on pairwise fusion for different models.
(a) ACM
(b) DBLP
(c) IMDB
Figure 8: The performance change among different datasets with the change of $\alpha$.
V-E Ablation Study
We study the impact of pairwise fusion in this part. We consider different candidates in the pairwise fusion process, aiming to find the best match that is capable of deriving high-quality features from the heterogeneous pairwise graph as the initial node embedding. We analyze the importance of pairwise fusion and implement different models introduced in Section V-A. Note that we applied the same hyperparameters of all these models according to the best performance reported in the respective papers and source codes. As shown in Figure 7, using PC-HGN in pairwise fusion enlarges the performance gap over other models in all datasets, with a performance gain of more than 5.4%-14.4% compared to using GAT, a performance gain of more than 6.6%-18.0% compared to using HGT, and an overall 11.5% performance gain on average compared to using other models in pairwise fusion. This result demonstrates the initial node embedding generated by PC-HGN helps our framework achieve the best results in all three datasets.
V-F Impact of Hyperparameter $\alpha$
As defined in Eq. 6, the proposed united loss is a linear combination of the reconstruction loss in dynamic heterogeneous hyperedges generation and the supervised loss for the downstream task. We further study the impact of $\alpha$ on the performance of LFH. In the experiment, the sensitivity of $\alpha$ is explored to control the level of impact caused by the reconstruction of each master when training the model. When the value of $\alpha$ is close to 1, it means that the model is optimized towards the reconstruction correctness. Fig. 8 reveals the change of F1 score along with $\alpha$. With $\alpha$ ranging in $\left[0.001,0.5\right]$, the performance first reaches to the top and then drops significantly, indicating that over-reliance on reconstruction loss aggravates the generalization ability of our model. We observe that F1 score reaches the highest when $\alpha=0.1$ in all three datasets. The result demonstrates that uniting reconstruction loss in the training process to some extent can help improve the model performance.
VI conclusion
In this paper, we propose a heterogenous hypergraph learning framework for representation learning. This framework first generates the high-quality initial node embedding using the designated pairwise fusion function, aiming to exploit the pairwise graph information at most. Afterwards, the multi-type hyperedges are constructed dynamically, forming the hypergraph together. The embedding is then updated iteratively through type-specific attention, aiming to encode the heterogeneous attributes into the embedding space. We conduct comprehensive experiments on three widely-used public datasets and comparison with eleven baseline methods to demonstrate the effectiveness of our proposed framework. The results and analysis reveal that the proposed framework can achieve new state-of-the-art performance on both node classification and link prediction tasks.